the sum of three numbers is 20.The second number is 4 times the first, and the sum of the first and third is 8. find the number
Answer by checkley71(8403)

x+y+z=20
y=4x
x+z=8 or z=8-x
x+(4x)+(8-x)=20
x+4x+8-x=20
4x=20-8
4x=12
x=12/4
x=3 answer
y=4*3=12 answer.
z=8-3=5 answer.
proof
3+12+5=20
20=20